# LV-UNet: A Lightweight and Vanilla Model for Medical Image Segmentation

### v1

@@ V-A1 ISIC2016
ISIC 2016 [11] is a dataset of dermoscopic images of skin lesions for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2016. In this paper, we used the lesion segmentation dataset, which contains original images and binary lesion masks. The training set has 900 pairs of original images and corresponding masks. In the testing set, there are 379 pairs of images and masks. The training set is split into a new training set and a validation set with the random state 42.
@@ V-A2 BUSI
Breast Ultrasound Images Dataset (Dataset BUSI) [12] is a dataset of breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, and corresponding masks. Unlike the setting in [16], we include all three classes in the dataset, and the test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
@@ V-A3 CVC-ClinicDB
CVC-ClinicDB [13] is a database of frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
@@ V-A4 CVC-ColonDB
The CVC-ColonDB dataset [14] includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
@@ V-A5 Kvasir-SEG
The Kvasir-SEG dataset [15] comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
Figure 3: Example segmentation results of different models on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvasir-SEG.
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [24]. The initial learning rate is 0.001, and a CosineAnnealingLR [25] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metrics in validation are $IOU$ of the lesions. The loss function used is a mixed loss combining binary cross entropy (BCE) loss and dice loss [26]:
@@
We also train, validate and test UNet, Attention UNet, UNet++, UNext, MLUNet, EGE-UNet, Rolling-UNet-M, U-KAN and MobileNetv3-UNet (UNet with MobileNetv3 as the backbone) for comparison. UNet uses double convolution layers in each stage. The dimensions of each stage in UNet, Attention UNet and UNet++ are [32, 64, 128, 256, 512], while the UNext, MLUNet, EGE-UNet, Rolling-UNet-M, and U-KAN use the original design in their papers. MobileNetv3-UNet adds the output of each encoder stage to the blocks in the decoder.
@@ V-C Results
The comparison experimental results can be seen in Table \@slowromancapi@, from which we can see LV-UNet shows competitive performances compared to the baselines. The deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and GFLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point-of-care scenarios. And some example segmentation results can be seen in Figure 3.
TABLE II: Parameters and GFLOPs of Different Models
Methods Parameters GFLOPs
UNet 7.8M 12.15
Attention-UNet 8.7M 16.79
UNet++ 9.2M 34.77
UNext 1.5M 0.58
MALUNet 0.2M 0.09
EGEUNet 53.4K 0.08
Rolling-UNet-M 7.1M 0.83
U-KAN 25.4M 10.5
MobileNetv3-Large-UNet 2.0M 0.51
LV-UNet 0.9M 0.22
LV-UNet (deployment) 0.5M 0.20
TABLE III: Experimental Results of Different Numebers of pre-trained Layers
@@
Combination 1 0.6308 0.7548 0.8531 0.9197 0.8063 0.8901
Combination 2 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
Combination 3 0.6332 0.7625 0.8324 0.9080 0.7971 0.8856
TABLE IV: Parameters and GFOPs of Different Numbers of pre-trained Layers
      Methods        Parameters       GFLOPs
      Combination 1       2.8M       0.34
      Combination 2       0.9M       0.22
      Combination 3       0.8M       0.17
@@ V-D Ablation Studies
The Number of Pre-trained Layers We explore the number of pre-trained layers of MobileNetv3-Large used in LV-UNet, and propose three possible combinations in the encoder:
@@
    Combination 1: From the initial convolutional stage to the inverted residual block $r14$ in MobileNetv3-Large and two encoder fusible blocks
@@
    Combination 2: From the initial convolutional stage to the inverted residual block $r9$ in MobileNetv3-Large and three encoder fusible blocks
@@
    Combination 3: From the initial convolutional stage to the iinverted residual block $r15$ in MobileNetv3-Large and four encoder fusible blocks
TABLE V: Experimental Results of Different Numebers of Series Informed Activation Functions
@@
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and GFLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting ( combination 2) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination 1, combination 2 reduces the model size and computational complexities significantly.
The Number of Series-Informed Activation Functions In experiments in [22], the increase of the number of Series-Informed Activation Functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. But the results can suggest that using an appropriate number of Series-Informed Activation Functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
@@
TABLE VII: Comparison Experimental Results of Different Nonlinearity Decreasing Methods for LV-UNet (Delopyment Mode)
@@
The Way of Skip-Connections We evaluated the two kinds of Skip-Connections methods: adding and concatenation. The adding operation in Skip-Connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies We also compared diffrent nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in [22] $a={e}/{E}$. The comparison results Table \@slowromancapvii@ show that performances are very close and our method achieves better results on more datasets.
@@ VI Conclusions
This paper proposes a lightweight and vanilla model named LV-UNet, which utilizes a pre-trained MobileNetv3-Large model and introduces fusible modules. The model can be trained using an improved deep training strategy and switched to deployment mode during inference, reducing both parameter count and computational load. Experiments are conducted on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvair-SEG datasets, achieving better performance compared to state-of-the-art and classic models. And the effect of the number of pre-trained layers, the number of Series-Informed Activation Functions, the way of skip-connections and deep training strategies are carefully studied. The research presents a general way to design a lightweight model for medical image segmentation: the combination of pre-trained models and fusible modules, which may provide some hints for further studies.
Acknowledgement

### v2

@@ V-A1 ISIC2016
ISIC 2016 [13] is a dataset of dermoscopic images of skin lesions for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2016. In this paper, we used the lesion segmentation dataset, which contains original images and binary lesion masks. The training set has 900 pairs of original images and corresponding masks. In the testing set, there are 379 pairs of images and masks. The training set is split into a new training set and a validation set with a ratio of 8:2.
@@ V-A2 BUSI
Breast Ultrasound Images Dataset (BUSI) [14] is a dataset of breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, and corresponding masks. Unlike the setting in [18], we include all three classes in the dataset, and the test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-A3 CVC-ClinicDB
CVC-ClinicDB [15] is a database of frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. The test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-A4 CVC-ColonDB
The CVC-ColonDB dataset [16] includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-A5 Kvasir-SEG
The Kvasir-SEG dataset [17] comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. The test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [26]. The initial learning rate is 0.001, and a CosineAnnealingLR [27] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metrics in validation are $IOU$ of the lesions. The loss function used is a mixed loss that combined binary cross-entropy (BCE) loss and dice loss [28]:
@@
We also train, validate and test UNet, Attention UNet, UNet++, UNeXt, MLUNet, EGE-UNet, U-KAN and MobileNetv3-UNet (UNet with MobileNetv3 as the backbone) for comparison. UNet uses double convolution layers in each stage. The dimensions of each stage in UNet, Attention UNet and UNet++ are 32, 64, 128, 256 and 512. UNeXt, MLUNet, EGE-UNet, and U-KAN use the original design in their papers. MobileNetv3-UNet adds the output of each encoder stage to the blocks in the decoder.
@@ V-C Results
The comparison experimental results can be seen in Table \@slowromancapi@, from which we can see LV-UNet shows competitive performances compared to the baselines. The deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and FLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point-of-care scenarios. Visualizations of some segmentation results can be seen in Figure 3.
TABLE II: Parameters and FLOPs of Different Models
Methods Parameters FLOPs
UNet 7.8M 12.15G
Attention-UNet 8.7M 16.79G
UNet++ 9.2M 34.77G
UNeXt 1.5M 0.58G
MALUNet 0.2M 0.09G
EGEUNet 53.4K 0.08G
U-KAN 25.4M 10.5G
MobileNetv3-Large-UNet 2.0M 0.51G
LV-UNet 0.9M 0.22G
LV-UNet (deployment) 0.5M 0.20G
TABLE III: Experimental Results of Different Combinations for pre-trained Layers
@@
Combination \@slowromancapi@ 0.6308 0.7548 0.8531 0.9197 0.8063 0.8901
Combination \@slowromancapii@ 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
Combination \@slowromancapiii@ 0.6332 0.7625 0.8324 0.9080 0.7971 0.8856
TABLE IV: Parameters and FLOPs of Different Combinations for pre-trained Layers
      Methods        Parameters       FLOPs
      Combination \@slowromancapi@       2.8M       0.34G
      Combination \@slowromancapii@       0.9M       0.22G
      Combination \@slowromancapiii@       0.8M       0.17G
@@ V-D Ablation Studies
The Number of Pre-trained Layers. We explore the number of pre-trained layers of MobileNetv3-Large used in LV-UNet, and propose three possible combinations in the encoder:
@@
    Combination \@slowromancapi@: From the initial convolutional stage to the inverted residual block $r14$ in MobileNetv3-Large and two fusible encoder blocks
@@
    Combination \@slowromancapii@: From the initial convolutional stage to the inverted residual block $r9$ in MobileNetv3-Large and three fusible encoder blocks
@@
    Combination \@slowromancapiii@: From the initial convolutional stage to the inverted residual block $r5$ in MobileNetv3-Large and four fusible encoder blocks
TABLE V: Experimental Results of Different Numbers of Series Informed Activation Functions
@@
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and FLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting (combination \@slowromancapii@) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination \@slowromancapi@, combination \@slowromancapii@ reduces the model size and computational complexities significantly.
The Number of Series-Informed Activation Functions. In experiments in [21], the increase of the number of series-informed activation functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. However, the results suggest that using an appropriate number of Series-Informed Activation Functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
@@
TABLE VII: Experimental Results of Different Nonlinearity Decreasing Methods for LV-UNet (Deployment Mode)
@@
The Way of Skip-Connections. We evaluated the two kinds of Skip-Connections methods: adding and concatenation. The adding operation in Skip-Connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies. We also compared different nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in  [21] $a={e}/{E}$. The comparison results in Table \@slowromancapvii@ show that performances are very close and our method achieves better results on more datasets.
@@ VI Conclusions
This paper proposes a lightweight and vanilla model named LV-UNet, which utilizes a pre-trained MobileNetv3-Large model and introduces fusible modules. The model can be trained using an improved deep training strategy and shifted to deployment mode during inference, reducing both the parameter count and computational load. Experiments are conducted on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvair-SEG datasets, achieving better performance compared to state-of-the-art and classic models. The effect of the number of pre-trained layers, the number of Series-Informed Activation Functions, the way of skip-connections and deep training strategies are carefully studied. The research presents a general way to design a lightweight model for medical image segmentation: the combination of pre-trained models and fusible modules, which may provide some hints for further studies.
Acknowledgment